# Multi-fidelity learning for interatomic potentials: Low-level forces and high-level energies are all you need

### v1

@@ 1 Introduction
Atomistic simulations, e.g., molecular dynamics (MD), find applications in many fields of science, including biology, materials science, and chemistry. The predictive capability of MD depends primarily on the accuracy of the forces acting on each atom. Historically, these forces have been calculated using either quantum mechanical (QM) methods, such as density functional theory (DFT), or classical mechanical methods, such as force fields (FFs), also referred to as interatomic potentials. Both QM and FF methods have limitations. Although QM calculations are very accurate, they are also extremely expensive, with a cost that increases rapidly with the number of electrons (typically to at least the third power). In comparison with QM, FFs are significantly less expensive and typically scale linearly with the number of atoms, rather than with the number of electrons. However, FFs are less accurate than QM because they use inflexible functional forms with relatively few parameters that are fit to limited amounts of data. Recently, machine learning interatomic potentials (MLIPs), and especially neural network based MLIPs, [1, 2] promise a “best-of-both-worlds” scenario — computational costs and scaling similar to classical FFs but accuracy similar to the QM method used to generate the training data. [3, 4]
The quality of an MLIP, both in terms of accuracy and robustness, is predicated on the training dataset, consisting of atomic positions with their corresponding energies and/or forces computed with the desired QM method. [5] Because neural networks are extremely flexible “universal function approximators”, developing reliable and transferable neural network MLIPs requires enormous amounts of training data. For this reason, many studies focus on developing large, accurate, and diverse QM datasets. [6, 7, 8, 9, 10]
@@
Another challenge limiting the impact of public training data is that most datasets are computed using relatively low-level QM methods, with DFT being the most common method of choice. [5] Although modern functionals can achieve reasonable accuracy given a large enough basis set, most QM datasets are computed with traditional functionals and moderate basis sets. Unfortunately, highly accurate QM methods (e.g., coupled-cluster, CC, with a complete basis set extrapolation) are orders of magnitude more computationally expensive. Consequently, the few datasets computed with high-level QM methods are much smaller in terms of both number of calculations and size of systems (i.e., number of atoms). [9]
A related challenge is that several QM codes cannot provide analytical forces for the most accurate QM methods. [12, 13, 14] Although most QM codes can compute numerical forces at any level of theory, these forces are more approximate and more expensive to compute as they require performing numerous energy calculations at slightly perturbed atomic configurations. For these reasons, although some recent efforts have expended significant computational resources to calculate large amounts of CC-level forces, [15] most datasets computed with high-level QM methods do not contain force data. [9, 16, 17, 18, 19] For example, while the ANI-1x dataset contains DFT-level energies and forces for $\approx$4.5M configurations of small organic molecules in vacuo, the ANI-1ccx dataset contains only CC-level energies for a subset of $\approx$500k configurations. [9]
@@
The abundance of incompatible low-accuracy datasets and the scarcity of high-accuracy forces requires more sophisticated training techniques that can simultaneously learn from multiple datasets. Figure 1 compares four such multi-dataset training paradigms, namely, transfer learning (TL), [18, 23, 25, 26] $\Delta$-learning ($\Delta$L), [17, 19, 27, 28] meta-learning ($\mu$L), [29] and multi-fidelity learning (MFL). [30, 31] All four methods present unique advantages and disadvantages when training to multiple datasets with different levels of theory. Both TL and $\Delta$L are powerful methods in the limit of only two datasets at different levels of theory, whereas $\mu$L and MFL are theoretically amenable to any number of levels of theory. Both TL and $\Delta$L consist of first training an MLIP on the large, low-accuracy dataset, followed by training to the smaller, high-accuracy dataset. By contrast, $\mu$L and MFL train to all datasets simultaneously, although $\mu$L also requires a subsequent fine-tuning step. While TL, $\Delta$L, and $\mu$L are two-step methods that result in separate MLIPs for each level of theory, MFL is a single-step method which trains a single MLIP that predicts multiple energies, with one output energy for each level of accuracy (i.e., level of fidelity). MFL is also often referred to as multi-level learning, multi-head learning, and multi-task learning (see Appendix for discussion of terminology). [32, 33, 34, 35]
@@
Figure 1: Comparison of different methods capable of training to multiple datasets at different levels of theory. Although meta-learning and multi-fidelity learning can utilize many datasets, for visual simplicity, only three levels of theory are depicted. Low-level data are represented with density functional theory (DFT). Intermediate-level data are represented with Møller-Plesset second-order perturbation theory (MP2). High-level data are represented with coupled-cluster theory (CC). While some variation exists in the literature as to the implementation of each method, the depictions in this figure are consistent with the work of Smith et al., Allen et al., and Jacobson et al. [23, 29, 30]
@@
Previous studies demonstrate that TL, $\Delta$L and $\mu$L achieve significant improvement when training to two datasets that overlap perfectly in configuration space, namely, the ANI-1x and ANI-1ccx datasets. [23, 29] Specifically, all three multi-dataset methods report approximately 20% higher accuracy compared to training solely to the ANI-1ccx CC-level energies. By contrast, previous MFL work did not see significant improvement when simultaneously training to ANI-1x and ANI-1ccx. [30] However, this study did not include the DFT-level atomic forces when performing MFL, only the DFT-level energies and CC-level energies were used for training. The justification for training only to energies was to avoid biasing the MLIP towards the ANI-1x DFT-level forces, since ANI-1ccx does not contain CC-level force data. Consequently, MFL with just DFT-level energies and CC-level energies achieved comparable accuracy to training solely to CC-level energies. Additional studies also tested the limits of MFL by training to several different levels of theory simultaneously. [30, 32, 36, 37] Learning from multiple datasets covering different regions of chemical space offers clear benefits for improved transferability, but does not necessarily lead to improved accuracy.
@@
The primary goal of this work is to demonstrate that MFL can result in an MLIP with extremely accurate high-level energies and forces, even when high-level forces are not available for training. Our hypothesis is that high-level forces can be learned implicitly with MFL by simultaneously training to low-level forces and high-level energies. To validate our hypothesis, we apply MFL to train an MLIP to the DFT-level energies and forces and CC-level energies for the $\approx$500k structures common to both the ANI-1x and ANI-1ccx datasets. These two levels of theory were chosen to enable comparison with previous studies that trained to the same datasets but utilized TL, [23] $\Delta$L, [23] $\mu$L, [29] and MFL without DFT-level forces. [30] For further validation, we repeat this MFL process but with two other intermediate-level QM methods in place of the CC-level energy data. For completeness, we briefly investigate any potential benefit of training with force data from more than two levels of theory.
@@ 2 Methods
All models in this study utilize the “Hierarchically Interacting Particle Neural Network” (HIP-NN) architecture, a state-of-the-art message-passing graph-based neural network MLIP. [38, 24] A detailed overview of HIP-NN is provided in the Appendix Section. Although previous studies have applied HIP-NN to predict multiple properties in addition to energy, such as atomic charge, [39] dipole, [40], bond orders [41], singlet-triplet excitation energies [42], and effective Hamiltonians[43, 44], HIP-NN has never been used to predict energy at multiple levels of theory. The HIP-NN architecture is fundamentally different from the “Accurate NeurAl networK engINe for Molecular Energies” (ANAKIN-ME, or ANI) feed forward neural network architecture used in previous MFL work. [2, 30] Although our primary conclusions regarding the performance of MFL do not depend on the MLIP architecture chosen, this section describes some MFL design decisions specific to HIP-NN.
@@ 2.1 Multi-fidelity learning with HIP-NN
Standard HIP-NN is already capable of predicting multiple outputs. Therefore, no modifications to the HIP-NN code were required to perform MFL. However, training HIP-NN to multiple levels of theory requires several important design decisions. The most important decision is which layers use shared parameters (i.e., parameters trained simultaneously to multiple levels of theory) and which layers use non-shared parameters (i.e., fidelity-specific parameters trained to a single level of theory). In this work, all interaction blocks use shared parameters while all linear output layers use fidelity-specific non-shared parameters (see Figure 2). Therefore, the weights and biases in Equation 4 and Equation 7 are shared parameters while the weights and biases in Equation 8 are the only non-shared parameters (see Appendix).
@@
where $t$ denotes the level of theory, $E$ stands for energy, $F$ stands for forces, $L_{E,t}$, $L_{F,t}$, and $L_{H,t}$ are the energy, force, and hierarchicality loss terms for the $t^{\rm th}$ level of theory, respectively, and $w_{E,t}$, $w_{F,t}$, $w_{H,t}$ are the weights for the energy, force, and hierarchicality loss terms for the $t^{\rm th}$ level of theory, respectively. The loss function for single-fidelity learning is obtained when the summation is over one level of theory. In this work, the weights for all contributions are equal, i.e., $w_{E,t}=1$, $w_{F,t}=1$, and $w_{H,t}=1$. However, force weights are zero $(w_{F,t}=0)$ if not training to forces from that level of theory. The $L_{H,t}$ term serves as a regularization term that ensures the MLIP obeys the principle of hierarchicality, namely, that the hierarchical energy contribution to the total energy decreases for increasing hierarchicality, i.e., $E^{n+1}<E^{n}$ (see Figure 2 and Appendix for details).
@@ 2.2.1 Training
Training is performed using a subset of the original ANI-1x dataset, which consists primarily of near-equilibrium (i.e., non-reactive) structures for stable organic molecules. [9] While the complete ANI-1x dataset contains $\approx$4.5M molecular structures, we utilize a subset of nearly 500k molecular structures that correspond to the same structures as the ANI-1ccx dataset. Both the complete ANI-1x dataset and the ANI-1ccx subset were generated in previous work through active learning. [45] Although the full ANI-1x dataset contains low-level energies and forces for all $\approx$4.5M structures, high-level energies are available only for this reduced subset. Thus, utilizing the reduced dataset of $\approx$500k structures helps avoid the issue of unequal data volumes between different QM methods. Previous studies have trained reliable MLIPs to both the complete ANI-1x dataset as well as significantly reduced subsets of ANI-1x. [24, 25, 46, 47, 48, 49]
The reduced ANI-1x dataset contains energies and forces computed with several QM methods. To simplify our analysis, we focus on only four levels of theory: $\omega$b97x/DZ, $\omega$b97x/TZ, MP2/TZ, and CCSD(T)/CBS. The two lowest fidelity levels ($\omega$b97x/DZ and $\omega$b97x/TZ) are based on DFT using the hybrid $\omega$b97x functional with either a double zeta (DZ) or triple zeta (TZ) basis set. [50, 51, 52] For brevity, we refer to these two DFT-based fidelity levels as DFT/DZ and DFT/TZ or simply as DZ and TZ, respectively. The second highest fidelity level corresponds to Møller-Plesset second-order perturbation theory with a triple zeta basis set (MP2/TZ). [53] The highest fidelity level utilizes an approximate coupled-cluster single and double excitation with perturbative triple excitation (CCSD(T)) with a complete basis set (CBS) extrapolation implementation. [9, 54] For brevity, we refer to the two highest fidelity levels as MP2 and CC, respectively. While energies are available for nearly all $\approx$500k structures across all four fidelity levels, forces are only available for the two DFT-based methods, namely, DZ and TZ.
@@ 2.2.2 Testing
Two types of test datasets are used to evaluate the accuracy of our trained models: in-sample and out-of-sample. In-sample test datasets are generated by randomly selecting 10% of the $\approx$500k configurations in the reduced ANI-1x dataset. These $\approx$50k configurations form a held-out dataset that is not included in training or validation. Each in-sample test dataset uses the same fidelity level as the training data.
Out-of-sample test datasets come from the GDB10to13 portion of the COMP6 dataset, [45] which contains approximately 48k conformations of larger molecules (10–13 heavy atoms) from the GDB-11 and GDB-13 databases. [55, 56, 57] For simplicity, we refer to this test dataset simply as COMP6, even though it only represents a subset of the entire COMP6 dataset. This test dataset primarily serves as an extensibility test. Forces in the COMP6 test dataset are computed using MP2/TZ, while conformer energy differences $(\Delta E)$ are computed with CCSD(T)/CBS. To allow for comparison with previous work and to focus on chemically relevant conformers, only conformers within 100 kcal/mol of the minimum-energy conformer are included in the analysis of COMP6 forces and $\Delta E$.

### v2

@@ 1 Introduction
Atomistic simulations, e.g., molecular dynamics (MD), find applications in many fields of science, including biology, materials science, and chemistry. The predictive capability of MD depends primarily on the accuracy of the forces acting on each atom. Historically, these forces have been calculated using either quantum mechanical (QM) methods, such as density functional theory (DFT), or classical mechanical methods, such as force fields (FFs), also referred to as interatomic potentials. Both QM and FF methods have limitations. Although QM calculations are very accurate, they are also extremely expensive, with a cost that increases rapidly with the number of electrons (typically to at least the third power). In comparison with QM, FFs are significantly less expensive and typically scale linearly with the number of atoms, rather than with the number of electrons. However, FFs are less accurate than QM because they use inflexible functional forms with relatively few parameters that are fit to limited amounts of data. Recently, machine learning interatomic potentials (MLIPs), [1] and especially neural network-based MLIPs, [2] promise a “best-of-both-worlds” scenario — computational costs and scaling similar to classical FFs but accuracy similar to the QM method used to generate the training data. [3]
While several different MLIP architectures exist, [4, 5, 6, 7, 8, 9, 10] the overall quality of an MLIP, both in terms of accuracy and robustness, [11] is primarily predicated on the training dataset. [12] By contrast to classical FFs, which are typically parameterized with limited amounts of experimental data, training datasets for MLIPs consist of atomic positions with their corresponding energies and/or forces computed with the desired QM method. Because neural networks are extremely flexible “universal function approximators”, developing reliable and transferable neural network MLIPs requires enormous amounts of training data. [13] For this reason, many studies focus on developing large and diverse QM datasets. [14, 15, 16, 17, 18, 19, 20, 21]
@@
Another challenge limiting the impact of public training data is that most datasets are computed using relatively low-level QM methods, with DFT being the most common method of choice. [22] Although modern functionals can achieve reasonable accuracy given a large enough basis set, most QM datasets are computed with traditional functionals and moderate basis sets. Unfortunately, highly accurate QM methods (e.g., coupled-cluster, CC, [24] with a complete basis set extrapolation) [25] are orders of magnitude more computationally expensive. Consequently, the few datasets computed with high-level QM methods are much smaller in terms of both number of calculations and size of systems (i.e., number of atoms). [17]
A related challenge is that several QM codes cannot provide analytical forces for the most accurate QM methods. [26, 27, 28] Although most QM codes can compute numerical forces at any level of theory, these forces are more approximate and more expensive to compute as they require performing numerous energy calculations at slightly perturbed atomic configurations. For these reasons, although some recent efforts have expended significant computational resources to calculate large amounts of CC-level forces, [29] most datasets computed with high-level QM methods do not contain force data. [17, 30, 31, 32, 33] For example, while the ANI-1x dataset contains DFT-level energies and forces for $\approx$4.5M configurations of small organic molecules in vacuo, the ANI-1ccx dataset contains only CC-level energies for a subset of $\approx$460k configurations. [17]
@@
The abundance of incompatible low-accuracy datasets and the scarcity of high-accuracy forces requires more sophisticated training techniques that can simultaneously learn from multiple datasets. Figure 1 compares four such multi-dataset training paradigms, namely, transfer learning (TL), [32, 37, 39, 40] $\Delta$-learning ($\Delta$L), [31, 33, 41, 42, 43] meta-learning ($\mu$L), [44] and multi-fidelity learning (MFL). [45, 46] All four methods present unique advantages and disadvantages when training to multiple datasets with different levels of theory. Both TL and $\Delta$L are powerful methods in the limit of only two datasets at different levels of theory, whereas $\mu$L and MFL are theoretically amenable to any number of levels of theory. Both TL and $\Delta$L consist of first training an MLIP on the large, low-accuracy dataset, followed by training to the smaller, high-accuracy dataset. By contrast, $\mu$L and MFL train to all datasets simultaneously, although $\mu$L also requires a subsequent fine-tuning step. Thus, $\Delta$L, and $\mu$L are two-step methods that result in separate MLIPs for each level of theory. By contrast, MFL is a single-step method which trains a single MLIP that predicts multiple energies, with one output energy for each level of accuracy (i.e., level of fidelity). MFL is also often referred to as multi-level learning, [47] multi-head learning, and multi-task learning [48, 49] (see Appendix for discussion of terminology).
@@
Figure 1: Comparison of different methods capable of training to multiple datasets at different levels of theory. Although meta-learning and multi-fidelity learning can utilize many datasets, for visual simplicity, only three levels of theory are depicted. Low-level data are represented with density functional theory (DFT). Intermediate-level data are represented with Møller-Plesset second-order perturbation theory (MP2). High-level data are represented with coupled-cluster theory (CC). While some variation exists in the literature as to the implementation of each method, the depictions in this figure are consistent with the TL and $\Delta$L work of Smith et al., [37] the $\mu$L work of Allen et al., [44] and the MFL work of Jacobson et al. [45]
@@
Previous studies demonstrate that TL, [37] $\Delta$L, [37] and $\mu$[44] achieve significant improvement when training to two datasets that overlap perfectly in configuration space, namely, the ANI-1x and ANI-1ccx datasets. Specifically, all three multi-dataset methods report approximately 20% higher accuracy compared to training solely to the ANI-1ccx CC-level energies. By contrast, previous MFL work did not see significant improvement when simultaneously training to ANI-1x and ANI-1ccx. [45] However, this study did not include the DFT-level atomic forces when performing MFL, only the DFT-level energies and CC-level energies were used for training. The justification for training only to energies was to avoid biasing the MLIP towards the ANI-1x DFT-level forces, since ANI-1ccx does not contain CC-level force data. Consequently, MFL with just DFT-level energies and CC-level energies achieved comparable accuracy to training solely to CC-level energies. Additional studies also tested the limits of MFL by training to several different levels of theory simultaneously. [45, 47, 50, 51] Learning from multiple datasets covering different regions of chemical space offers clear benefits for improved transferability, but does not necessarily lead to improved accuracy.
@@
The primary goal of this work is to demonstrate that MFL can result in an MLIP with extremely accurate high-level energies and forces, even when high-level forces are not available for training. Our hypothesis is that high-level forces can be learned implicitly with MFL by simultaneously training to low-level forces and high-level energies. To validate our hypothesis, we apply MFL to train an MLIP to the DFT-level energies and forces and CC-level energies for the $\approx$460k structures common to both the ANI-1x and ANI-1ccx datasets. These two levels of theory were chosen to enable comparison with previous studies that trained to the same datasets but utilized TL, [37] $\Delta$L, [37] $\mu$L, [44] and MFL without DFT-level forces. [45] For further validation, we repeat this MFL process but with two other intermediate-level QM methods in place of the CC-level energy data. For completeness, we briefly investigate any potential benefit of training with force data from more than two levels of theory.
@@ 2 Methods
All models in this study utilize the “Hierarchically Interacting Particle Neural Network” (HIP-NN) architecture, a state-of-the-art message-passing graph-based neural network MLIP. [52] A detailed overview of HIP-NN is provided in the Appendix Section. The HIP-NN architecture is fundamentally different from the “Accurate NeurAl networK engINe for Molecular Energies” (ANAKIN-ME, or ANI) [5] feed forward neural network architecture used in previous MFL work. [45] Although our primary conclusions regarding the performance of MFL do not depend on the MLIP architecture chosen, this section describes some MFL design decisions specific to HIP-NN.
@@ 2.1 Multi-fidelity learning with HIP-NN
Although HIP-NN has never been used previously to predict energy at multiple levels of theory, standard HIP-NN is already capable of predicting multiple outputs. Therefore, no modifications to the HIP-NN code were required to perform MFL. However, training HIP-NN to multiple levels of theory requires several important design decisions. The most important decision is which layers use shared parameters (i.e., parameters trained simultaneously to multiple levels of theory) and which layers use non-shared parameters (i.e., fidelity-specific parameters trained to a single level of theory). In this work, all interaction blocks use shared parameters while all linear output layers use fidelity-specific non-shared parameters (see Figure 2). Therefore, the weights and biases in Equation 4 and Equation 7 are shared parameters while the weights and biases in Equation 8 are the only non-shared parameters (see Appendix).
@@
where $t$ denotes the level of theory, $E$ stands for energy, $F$ stands for forces, $L_{E,t}$, $L_{F,t}$, and $L_{H,t}$ are the energy, force, and hierarchicality loss terms for the $t^{\rm th}$ level of theory, respectively, and $w_{E,t}$, $w_{F,t}$, $w_{H,t}$ are the weights for the energy, force, and hierarchicality loss terms for the $t^{\rm th}$ level of theory, respectively. The loss function for single-fidelity learning is obtained when the summation is over one level of theory. In this work, the weights for all contributions are equal, i.e., $w_{E,t}=1$, $w_{F,t}=1$, and $w_{H,t}=1$. However, force weights are zero $(w_{F,t}=0)$ if not training to forces from that level of theory. Several other studies also use an equal energy-to-force weight ratio when training diverse MLIP architectures to the ANI-1x dataset. [53, 54, 55, 56] Although the present work utilizes constant weights throughout training, future work could investigate whether an adaptive loss weighting approach is beneficial for MFL. [57] Specifically, the low-level forces would be assigned greater weights at early epochs while the high-level energies would be assigned greater weights at later epochs, similar to recent work with knowledge distillation. [58]
The $L_{H,t}$ term serves as a regularization term that ensures the MLIP obeys the principle of hierarchicality, namely, that the hierarchical energy contribution to the total energy decreases for increasing hierarchicality, i.e., $E^{n+1}<E^{n}$ (see Figure 2 and Appendix for details).
@@ 2.2.1 Training
Training is performed using a subset of the original ANI-1x dataset, which consists primarily of near-equilibrium (i.e., non-reactive) structures for stable organic molecules. [17] While the complete ANI-1x dataset contains $\approx$4.5M molecular structures, we utilize a subset of nearly 460k molecular structures that correspond to the same structures as the ANI-1ccx dataset. Both the complete ANI-1x dataset and the ANI-1ccx subset were generated in previous work through active learning. [59] Although the full ANI-1x dataset contains low-level energies and forces for all $\approx$4.5M structures, high-level energies are available only for this reduced subset. Thus, utilizing the reduced dataset of $\approx$460k structures helps avoid the issue of unequal data volumes between different QM methods. Previous studies have trained reliable MLIPs for diverse architectures to both the complete ANI-1x dataset as well as significantly reduced subsets of ANI-1x. [38, 39, 53, 54, 55, 56, 60, 61]
The reduced ANI-1x dataset contains energies and forces computed with several QM methods. To simplify our analysis, we focus on only four levels of theory: DFT/DZ, DFT/TZ, MP2/TZ, and CCSD(T)*/CBS. The two lowest fidelity levels (DFT/DZ and DFT/TZ) are based on the hybrid $\omega$b97x functional [62] with either the 6-31G* double zeta (DZ) basis set [63] or the def2-TZVPP triple zeta (TZ) basis set. [64] For brevity, we refer to these two DFT-based fidelity levels simply as DZ and TZ, respectively. The second highest fidelity level (MP2/TZ) corresponds to Møller-Plesset second-order perturbation theory (MP2) [65] with the cc-pVTZ triple zeta basis set. [66, 67] The highest fidelity level (CCSD(T)*/CBS) utilizes an approximate coupled-cluster (CC) single and double excitation with perturbative triple excitation [68] with a complete basis set (CBS) extrapolation implementation [69, 70] (see Smith et al. [17] for details). For brevity, we refer to the two highest fidelity levels as MP2 and CC, respectively. While energies are available for nearly all $\approx$460k structures across all four fidelity levels, forces are only available for the DFT/DZ and DFT/TZ methods.
@@ 2.2.2 Testing
Two types of test datasets are used to evaluate the accuracy of our trained models: in-sample and out-of-sample. In-sample test datasets are generated by randomly selecting 10% of the $\approx$460k configurations in the reduced ANI-1x dataset. These $\approx$50k configurations form a held-out dataset that is not included in training or validation. Each in-sample test dataset uses the same fidelity level as the training data.
Out-of-sample test datasets come from the GDB10to13 portion of the COMP6 dataset, [59] which contains approximately 48k conformations of larger molecules (10–13 heavy atoms) from the GDB-11 database [71, 72] and the GDB-13 database. [73] The GDB10to13 test dataset primarily serves as an extensibility test. Forces in the GDB10to13 test dataset are computed using MP2/TZ, while conformer energy differences $(\Delta E_{\rm conf})$ are computed with CCSD(T)*/CBS. To allow for comparison with previous work and to focus on chemically relevant conformers, only conformers within 100 kcal/mol of the minimum-energy conformer are included in the analysis of GDB10to13 forces and $\Delta E_{\rm conf}$.
An additional out-of-sample test dataset contains torsion scan energy differences $(\Delta E_{\rm tors})$ for small drug-like molecules. This test dataset contains the same geometries as Sellers et al. [74] that were previously optimized with MP2 and the 6-311+G** triple zeta basis set. [75] However, the energies in our torsion scan test dataset were previously recomputed using the same CCSD(T)*/CBS approach as the ANI-1ccx dataset. We also test our models against the CCSD(T)/CBS energies reported by Sellers et al. (see Supporting Information), to allow for a fair comparison with other studies. [37, 45, 58, 76] To further test the robustness of our models, we also perform relaxed torsion scans, wherein the positions of all atoms are optimized with a given model subject to the constraint of a fixed dihedral angle (see Supporting Information).
Performance for each model is measured by the mean absolute error (MAE) and/or root-mean-square error (RMSE) of energies and atomic forces. The RMSE and MAE are reported as an average value with a corresponding 95% confidence interval from an ensemble of eight MLIPs. Specifically, the RMSE and MAE are first computed for each individual ensemble member. Subsequently, the mean and standard deviation are computed from the eight different RMSE and MAE values. Consequently, the average RMSEs and MAEs reported are not based on ensemble-averaged energies and forces. Averaging the energy and forces over an ensemble of MLIPs smooths the potential energy surface and, thereby, significantly reduces the errors. However, by not averaging the predictions prior to computing the errors, it is possible to quantify the ensemble uncertainty in the errors themselves and, thus, determine whether two models are statistically different. All errors reported in this work correspond to test errors, i.e., the errors computed on either the held-out dataset or the GDB10to13 testset. No training or validation errors are reported.